OpenStudy (anonymous):

What does... 2a - 2b - a^2 +b^2 simplify to?? (Using factoring)

OpenStudy (anonymous):

The entire problem is: \[\frac{ 2a + 2b - a^{2} + b^2 }{ 9a^2 - 6ab +b^2 } \times \frac{ 9a^2 - b^2 }{ 3a^2 -2ab-b^2 }\]

OpenStudy (anonymous):

the first "+" should be a "-"